OpenStudy (anonymous):

How much salt is in a mixture containing 10 gallons of an 11% salt solution and 6 gallons of a 7% salt solution?

OpenStudy (anonymous):

- 10 gallons with 11% salt: there is 0.11 to 1.0 salt, so there are 1.1 gallons salt - 6 gallons with 7% salt: there is 0.07 to 1.0 salt, and 1.0 ~ 6 gallons: 6 x 0.07 or 0.42 gallons of salt - in total: 1.1 gallons + 0.42 gallons = 1.52 gallons of salt.

OpenStudy (anonymous):

Thank you so much. WHat would the percentage of pure water contained in this same mixture be?

OpenStudy (anonymous):

first find the % of the salt: we have 1.52 gallons salt in 10+6 gallons of water or 1.52 gall in 16 gall to get a percentage we must compare with 100, not 16

OpenStudy (anonymous):

then just subtract 1.52 from 16 gal?

OpenStudy (anonymous):

we must make the denominator 100 while conserving the ratio possible way achieving this: 1,52 * 100/16 ------------- 16 * 100/16

OpenStudy (anonymous):

want 100 in the denominator, so I multiply with 100 however there was already a 16?! so I also divide 16 to get rid of it then only 100 is left in the denominator, like should be for % we have to apply the same operations to the numerator: if we do this, the ratio did not change, just the sizes

OpenStudy (anonymous):

Karen is working in the chemistry lab and needs a solution that is 8% hydrochloric acid. She has two solutions that she can mix together to make the solution of the required strength. The first solution is 2% hydrochloric acid, and the other is 11% hydrochloric acid. How many kilograms of each solution should she mix to produce 3 kilograms of the solution of the required strength?

OpenStudy (anonymous):

how would I do this

OpenStudy (anonymous):

do you know the percentage of clear water for the prev one?

OpenStudy (anonymous):

1,52 * 100/16 ------------- 16 * 100/16 that's percentage of salt in new solution. percentage of pure water is just 100% - above % because they ask for PERCENTAGE and not parts of 16 so you can't just subtract 1.52 from 16 g because that's not a real %

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

I got 9,5% salt in the solution so it's 90,5% pure water in the solution

OpenStudy (anonymous):

Karen is working in the chemistry lab and needs a solution that is 8% hydrochloric acid. She has two solutions that she can mix together to make the solution of the required strength. The first solution is 2% hydrochloric acid, and the other is 11% hydrochloric acid. How many kilograms of each solution should she mix to produce 3 kilograms of the solution of the required strength?

OpenStudy (anonymous):

you could make an equation for that. is that a good way or are you supposed to solve this question with a different technique?

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

I really appreciate your help.

OpenStudy (anonymous):

11 = x*(0.02) + y*(0.11) and also x + y = 3 the reason to have 2 equations is: there could be more than one way to get 8% hydrochloric acid, for example she could put double 11% and also double the 2% solution, it would give the same concentration.

OpenStudy (anonymous):

and that also makes things difficult....

OpenStudy (anonymous):

no doubt

OpenStudy (anonymous):

anyway it's possible to solve this SYSTEM OF EQUATIONS

OpenStudy (anonymous):

but since it is all hypochloric acid, would I need x & y

OpenStudy (anonymous):

x & y are the solutions of 11% and 2% basically like two plastic containers, one of them has 11% and the other 2% even though they all contain the same stuff, it's important to follow the rule that you e.g. take 2 bottles of the one and 1 bottle of the other

OpenStudy (anonymous):

otherwise it is hydrochloric solution however the concentration is wrong

OpenStudy (anonymous):

SO how then would you solve this problem?

OpenStudy (anonymous):

you can solve a system of equations by 1) subsitution 2) elimination

OpenStudy (anonymous):

which one do you prefer? I don't know which one is better in this case they both work

OpenStudy (anonymous):

would I set it up like this? .02x + .11(3 - x) = .24

OpenStudy (anonymous):

yes, this is solving by substitution and that's a correct setup

OpenStudy (anonymous):

where did you get .24 ?

OpenStudy (anonymous):

from the 8% new solution, I took 3 kilograms x .08 = .24

OpenStudy (anonymous):

so would the correct answer be 1 kilogram of the 8% and 2 kilograms of the 11%

OpenStudy (anonymous):

we're already considering the 3kg in the y=3-x I believe

OpenStudy (anonymous):

we can check if that's correct: concentration (1*8 + 2*11 )/3 = (22+8)/3 = 10% hydrochl.

OpenStudy (anonymous):

this equation was wrong, 11 = x*(0.02) + y*(0.11) and also it should be 0.08 = (0.02x + 0.11y)/(x+y)

OpenStudy (anonymous):

so in the numerator we count the amount of salt that flows into the mix and in the denominator we keep track of the total kg to be able to compute the %

OpenStudy (anonymous):

and the % should be 8% or 8/100 = 0.08

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok so we have these two equations that must both be satisfied: 1.) 0.08 = (0.02x + 0.11y)/(x+y) 2.) x + y = 3

OpenStudy (anonymous):

we can substitute y = 3-x in the first equation: \[0.08 = \frac{(0.02x + 0.11[3-x])}{ (x+[3-x])}\]

OpenStudy (anonymous):

and expand the substitute: \[0.08 = \frac{(0.02x + 0.33-0.11x)}{ (x+3-x)}\] we can combine like terms: \[0.08 = \frac{(-0.09x + 0.33)}{ (3)}\]

OpenStudy (anonymous):

and now, algebra - multiply both sides by 3: \[0.24 = -0.09x + 0.33\] and then move x to the left, move 0.24 to the right: \[0.09x =+ 0.33 -0.24\] combine the constants \[0.09x =+ 0.09\]

OpenStudy (anonymous):

so x=1

OpenStudy (anonymous):

yes and the fact we got such a whole number indicates it might be correct

OpenStudy (anonymous):

x+y = 3 y = 3-x y= 3-1 =2

OpenStudy (anonymous):

you already got that answer above maybe while checking it I confused 11% and 2% ??

OpenStudy (anonymous):

x=2%, and y=11% so if x=1 and y=2 then the concentration is: (2% + 2*11%)/3 = (2+22 )/3 = 24/3 = 8%

OpenStudy (anonymous):

ok you already had the right solution before all of the fractions I just didn't check it correctly at least now we have peer - reviewed the answer!!

OpenStudy (anonymous):

she must put 1kg of the 2% solution 2 kg of the 11% solution to get 3 kg of 8% solution

OpenStudy (anonymous):

One more please

OpenStudy (anonymous):

Susie is making a mixture of acetic acid and water, which forms vinegar. If she has a mixture of 100 ml of 2% acetic acid, how many ml of pure acetic acid should she add to produce the required 8% acetic acid vinegar? (Round to the nearest tenth.)

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

100 * 0.2 + x * 1.0 = 0.8/(100+x)

OpenStudy (anonymous):

great

OpenStudy (anonymous):

so x = .16 ??

OpenStudy (anonymous):

then solve for x 100 * 0.2 + x * 1.0 = 0.8/(100+x) (20 + x) *(100+x) = 0.8 2000 + 20x + 100 x + x^2 = 0.8 x^2 +120x + 2000 = 0.8 x^2 + 120x + 1999.2 = 0

OpenStudy (anonymous):

you lost me there... so the answer is 0 ??

OpenStudy (anonymous):

the eq was wrong (100 * 0.2 + x * 1.0) /(100+x) = 0.8

OpenStudy (anonymous):

(100 * 0.2 + x * 1.0) = 0.8 * (100+x) 20+x = 80+0.8x

OpenStudy (anonymous):

20+0.2x = 80 0.2x = 60 2 x = 600 x = 300

OpenStudy (anonymous):

thats clearly wrong Susie is making a mixture of acetic acid and water, which forms vinegar. If she has a mixture of 100 ml of 2% acetic acid, how many ml of pure acetic acid should she add to produce the required 8% acetic acid vinegar? (Round to the nearest tenth.)

OpenStudy (anonymous):

do you come up with .4

OpenStudy (anonymous):

I see why, I used 0.2 for the 2% acid I TREATED IT AS 20% acid

OpenStudy (anonymous):

here with adjustment: (100 * 0.02 + x * 1.0) /(100+x) = 0.08

OpenStudy (anonymous):

that is what i was thinking so did you come up with .16

OpenStudy (anonymous):

you got a good idea about this stuff I didn't come up with anything yet: (100 * 0.02 + x * 1.0) = 0.08 * (100+x) 2 + x = 8+0.08x

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

2 + x = 8+0.08x |subtract 0.08x from both sides 2 + 0.92 x = 8 0.92 x = 6 |(x100) 92 x = 600 x = 600/92

OpenStudy (anonymous):

beleive it or not i am really trying to work it out, just working with the word problems throws me off

OpenStudy (anonymous):

so I get 6,52 ml of pure acid are sufficient to elevate 100ml of 2% upto 8%

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

I guess thats plausable because we wanna add like 10% of acidity so we add like a tenth of pure acid....

OpenStudy (anonymous):

this will make an old man out of me

OpenStudy (anonymous):

so we are not looking at 6.5

OpenStudy (anonymous):

we are because the solution already had 2% and the real values are off by some small percentage due to the pure acid itself increasing volume

OpenStudy (anonymous):

like if we had 100ml of 0% and we add 10ml of 100% then we have 110 ml and it's slightly less than 10%

OpenStudy (anonymous):

God bless you!! THanks for the help

OpenStudy (anonymous):

thats why we need 6,5 ml and not just 6 ml to get 2% upto 8%